# Bridge indices of spatial graphs and diagram colorings

### v1

@@ Abstract
We extend the Wirtinger number of links, an invariant originally defined by Blair et al. in terms of extending initial colorings of some strands of a diagram to the entire diagram, to spatial graphs. We prove that the Wirtinger number equals the bridge index of spatial graphs, and we implement an algorithm in Python which gives a more efficient way to estimate upper bounds of bridge indices. Combined with lower bounds from diagram colorings by elements from certain algebraic structures, we obtain exact bridge indices for a large family of spatial graphs.
@@ 1 Introduction
The bridge index has been a useful complexity measure of knots and links since the 1950s, when Schubert used it to show that a knot has only finitely many companions [schubert1954numerische]. Since then, the bridge index has also been used to estimate several geometric quantities such as distortion [pardon2011distortion, blair2020distortion], lattice stick number [adams2012stick], and total curvature [milnor1950total]. It is conjectured that the bridge index equals an algebraic quantity called the meridional rank; a counterexample has not yet been discovered (see Problem 1.11 in Kirby’s list [kirby1997problems]). Blair et al. verified that this conjecture holds for a large family of links by rephrasing the bridge index algorithmically [blair2020wirtinger, baader2021coxeter].
Knots and links can be considered as a proper subclass of spatial graphs, which encompass embeddings of 1-complexes. These graphs embedded in 3-space are objects of interest to chemists, as atoms of a molecule and their chemical bonds can be modeled as vertices and edges of a spatial graph, respectively [castle2008ravels]. Additionally, conclusions about abstract graphs can be made by considering how they are spatially embedded in space. For example, the complete graphs $K_{6}$ and $K_{7}$ fall into two distinct levels in a filtration of abstract graphs formed by considering the minimum number of links in any embedding of the graph: any embedding of $K_{6}$ (resp. $K_{7}$) in 3-space contains at least one link (resp. 21 links) [fleming2009counting].
Various authors have generalized the bridge index to spatial graphs, and some have shown that their version still approximates curvature as the bridge index of knots does [soma2002curvature, gulliver2012total]. There are some properties however that do not carry over to the more general setting. For instance, any knot with bridge index exactly two is prime [schubert1956knoten], but this is no longer true for $\Theta$-graphs [motohashi20002]. For non-prime spatial graphs, there exist estimates for the bridge indices in terms of the number of summands [taylor2021tunnel].
Since a random spatial graph encountered in nature usually does not have a simple diagram, it is beneficial to generalize the method of [blair2020wirtinger] to systematically compute the bridge indices of spatial graphs from inputs that are computer-friendly. Our procedure can be summarized as follows. Given a spatial graph diagram, we remove vertices of degree at least three to get diagrams with free ends. We then pick a subset of strands to color and define a move that extends a color to an additional strand provided some conditions are met at a crossing. Roughly, the Wirtinger number is the minimum number of initial colorings that extend to the entire diagram by the move, minimized over all diagrams (see Section 2.3 for a formal definition). This quantity turns out to equal the bridge index, which we prove in Section 3.
@@
We use our main result, which we implement as an algorithm in Python, to provide upper estimates for the bridge index of a variety of spatial graphs. In some cases we combine our techniques with algebraic lower bounds, which we detail in Section 4, to compute exact values. We use our methods to give several short proofs in Section 5 that the bridge index of spatial graphs can be arbitrarily large (see Corollary 5.9, Corollary 5.13, and Corollary 5.18).
Corollary 1.2.
The bridge index of a spatial graph can be arbitrarily large.
Throughout Section 5 we provide many example computations. In particular, spatial graph whose vertices all have degree four can be obtained from link diagrams by turning some crossings into vertices, or fusing maxima and minima. We can compute exact bridge indices for spatial graphs obtained from generalized Montesinos links by fusing some maxima and minima together (see Example 5.17). In some cases, these fusing operations can still produce interesting almost unknotted graphs. Our Python code can be found at https://github.com/hanhv/graph-wirt.
@@ Acknowledgments
Some of this work was conducted when the first author visited the other two at Arizona State University, and she thanks the mathematics department for their hospitality. We would like to thank Julien Paupert for support and helpful advice. SB was supported by the NSF Postdoctoral Research Fellowship DMS-2303143.
@@ 2.1 Spatial graphs
A graph $G$ is a set $(V,E)$ of vertices and edges, where $G=E\subset V\times V$. A spatial graph is an embedding of a graph $G$ in $S^{3}$ taking vertices to points in $S^{3}$ and taking edges $(u,v)$ to an arc whose endpoints are images of $u$ and $v$. We say a spatial graph is trivial if it can be isotoped to lie in the plane.
@@
A straightforward way to construct a spatial graph is to start with a link, and then add some edges. This way, the complexity of the knotted graph is obtained from the original link. This motivates the study of the following class of spatial graphs. A spatial graph is almost unknotted if all of its proper subgraphs are trivial. For $\Theta$-graphs, this means each constituent knot, that is, the knot obtained by deleting one of the three edges, is the unknot.
@@ 2.2 Bridge splittings
For our definition of bridge splitting, we follow Taylor and Tomova [taylor2013c]. We call a properly embedded graph $\tau$ in a 3-ball $B$ a trivial tangle if the following two conditions are satisfied:
  1. (1)
    each connected component of $\tau$ is either a bridge arc, which is an edge with both endpoints on $\partial B$, or a pod, which is a graph with a single vertex in the interior of $B$ and with all degree one vertices contained in $\partial B$, and
  2. (2)
    all components of $\tau$ are simultaneously boundary parallel into $\partial B$. This means that for each component $t$ of $\tau$ there is a disk $D$ containing $t$ with $\partial D$ inessential in $\partial B$.
The degree of a pod is defined to be the degree of the single vertex in the interior of $B$. By convention, we assume pods have degree $\geq 3$. (Degree two vertices are treated as parts of bridge arcs.) See Figure 1 for an example of a trivial tangle. We will glue these trivial tangles along these degree one vertices to form closed spatial graphs in $S^{3}$, and in general, we think of the resulting closed spatial graphs as no longer having vertices where we glued.
Figure 1: A trivial tangle, where we have displayed pods with degrees two and three, but in general we allow pods of any degrees. We did not include vertices where the tangle intersects $\partial B$, since when we glue two trivial tangles together along these vertices, we think of the resulting closed spatial graph as no longer having vertices where we glued.
A bridge splitting of a spatial graph $G$ in the $3$-sphere $S^{3}$ is a decomposition into two trivial tangles $(S^{3},G)=(B_{1},\tau_{1})\cup(B_{2},\tau_{2})$. We sometimes refer to the pods in $(B_{1},\tau_{1})$ as upper pods and the pods in $(B_{2},\tau_{2})$ as lower pods. For bridge splittings of links in $S^{3}$, each tangle has the same number of maxima, and we call the minimum number of maxima over all possible bridge splittings the bridge index. Note however that in general the tangles in a bridge splitting of a spatial graph may have a different number of maxima. For this reason there exist multiple definitions of bridge index for spatial graphs in the literature. Our definition will be the following.
@@
The bridge index of a particular bridge splitting $(S^{3},G)=(B_{1},\tau_{1})\cup(B_{2},\tau_{2})$ of a spatial graph $G$ is the minimum number of connected components $|\tau_{1}|$. The bridge index of a spatial graph $G$, denoted $\beta(G)$, is the minimum bridge index over all bridge splittings for $G$.
When it is clear from context, we may abuse notation and use $\beta(G)$ to denote the bridge index of a particular bridge splitting of a graph as well. In the case of $\Theta_{n}$-graphs, the discrepancy around defining the bridge index disappears, as the two trivial tangles in a bridge splitting of a $\Theta_{n}$-graph will always have the same number of maxima, provided that we place one vertex in each tangle.
Remark 2.2.
When $G$ is a $\Theta_{n}$-graph and we require that each $\tau_{i}$ contains only one vertex of $G$, Definition 2.1 reduces to the definitions given in [goda1997bridge, motohashi20002, tsuno2003bridge].
Remark 2.3.
Once the number of components of one of the trivial tangles $|\tau_{1}|$ is known, one can calculate the number of components of the other trivial tangle $|\tau_{2}|$ in a bridge splitting as follows. Let $\chi(G)$ be the number of vertices of $G$ minus the number of edges of $G$. To keep track of the local maxima, we put a vertex of degree two at each local maximum which is a bridge arc. At the moment, the number of vertices is $|\tau_{1}|+|\tau_{2}|.$ The number of edges can then be determined by examining the sum of the degrees of the vertices in $\tau_{1}$: $\sum_{1\leq i\leq|\tau_{1}|}d_{i}$. It follows that
@@ 2.3 Wirtinger number
Given a spatial graph $G$ in $S^{3}$, a diagram $D$ for $G$ is a generic immersion of $G$ into the plane with each crossing labeled as an overcrossing or an undercrossing. Our new formulation of the bridge index starts with the formation a modified diagram from the original spatial graph diagram.
@@
Let $D$ be a diagram for a spatial graph. A truncated diagram is the result of removing vertices of degree $d>2$ from $D$. We refer to the remaining pieces of the graph near removed vertices as free ends. A strand of a truncated diagram is one of the following:
  1. (1)
    an arc connecting an undercrossing to a free end, or
  2. (2)
    an arc connecting two undercrossings.
We now describe a process for coloring a truncated diagram. First choose $k$ distinct colors, and color a subset of strands of $D$ using these $k$ colors, where the coloring does not repeat unless a color is assigned to strands that came from removing a degree $d>2$ vertex. In the latter case, we use the same color. These $k$ initial colors are called seeds. Sometimes we will also refer to the strands which are colored the initial colors as seeds (where strands coming from the removal of a degree $d>2$ vertex count as one seed, as they are all one color).
Then continue to color any uncolored strands using the following coloring move, which extends the coloring of a colored arc which passes under another colored arc to the other arc passing under the overstrand. See Figure 2. Note that the coloring move is valid both when the overstrand matches and does not match the color of the understrand; all that matters is that the overstrand has been colored. Also, while strands which came from the removal of a degree $d>2$ vertex must be colored the same color if they are included as a seed, such strands in general can end up being colored with different colors if not included as a seed.
@@
If a truncated diagram for a spatial graph can be completely colored by starting with $k$ seeds and extending these colors using the coloring move, we call the corresponding non-truncated diagram $k$-Wirtinger colorable. See Figure 3 for an example of a $3$-Wirtinger colorable diagram for a spatial $\Theta$-graph. It is an exercise to check that this diagram is not 2-Wirtinger colorable. In fact, our Python code checks for all possible combinations of strands given a fixed number of colors.
Figure 3: This diagram is 3-Wirtinger colorable.
In a $k$-Wirtinger colored diagram at the final stage, there are certain types of crossings that play an important role in our results. A crossing is called a multi-colored crossing if the corresponding understrands received different colors.
@@
The Wirtinger number of a spatial graph diagram $D$, denoted $\omega(D)$, is the minimal value of $k$ for which the diagram is $k$-Wirtinger colorable; in other words, the smallest number of seeds required to color the entire diagram. The Wirtinger number of a spatial graph $G$, denoted $\omega(G)$, is the minimal Wirtinger number over all diagrams for $G$.
Remark 2.6.
From the proof of our main theorem, we will see that the number $|\tau_{2}|$ can also be determined from a $k$-Wirtinger colored diagram as the number of multi-colored crossings plus the number of non-seed pods.
2.4 Braids
An $n$-string braid is a set of $n$ strings in $\mathbb{R}^{3}$ with one endpoint of each string attached to a horizontal plane $z=b$, and the other endpoint attached to another horizontal plane $z=a$, where $a<b$, with the additional requirement that as we move along each string from the top plane to the bottom plane, the string heads downwards and never backtracks. An observation we will frequently use is that as we travel along a spatial graph, we turn around to travel in the opposite direction at the the local extrema.
Observation 2.1.
Suppose that for a bridge-split spatial graph the upper pods and local maxima are at the same level. Suppose also that at a lower height, the lower pods and local minima are at the same level. Then removing small open neighborhoods of the vertices at the pods and the local maxima and minima results in a braid.

### v2

@@ Abstract
We extend the Wirtinger number of links, an invariant originally defined by Blair, Kjuchukova, Velazquez, and Villanueva in terms of extending initial colorings of some strands of a diagram to the entire diagram, to spatial graphs. We prove that the Wirtinger number equals the bridge index of spatial graphs, and we implement an algorithm in Python which gives a more efficient way to estimate upper bounds of bridge indices. Combined with lower bounds from diagram colorings by elements from certain algebraic structures and clasping techniques, we obtain exact bridge indices for a large family of almost unknotted spatial graphs. We also show that for every possible negative Euler characteristic, there exist almost unknotted graphs of arbitrarily large bridge index.
@@ 1 Introduction
The bridge index has been a useful complexity measure of knots and links since the 1950s, when Schubert used it to show that a knot has only finitely many companions [Sch54]. The bridge index has also been used to estimate several geometric quantities such as distortion [Par11, Bla+20], lattice stick number [Ada+12], and total curvature [Mil50]. It is conjectured that the bridge index equals an algebraic quantity called the meridional rank; a counterexample has not yet been discovered (see Problem 1.11 in Kirby’s list [Kir97]). Baader, Blair, and Kjuchukova verified that this conjecture holds for a large family of links by rephrasing the bridge index algorithmically [BBK21], building off of work by Blair, Kjuchukova, Velazquez, and Villanueva [Bla+20a].
Knots and links can be considered as a proper subclass of spatial graphs, which encompass embeddings of 1-complexes. These graphs embedded in 3-space are objects of interest to chemists, as atoms of a molecule and their chemical bonds can be modeled as vertices and edges of a spatial graph, respectively [CEH08]. Additionally, conclusions about abstract graphs can be made by considering how they are spatially embedded in space. For example, the complete graphs $K_{6}$ and $K_{7}$ fall into two distinct levels in a filtration of abstract graphs formed by considering the minimum number of links in any embedding of the graph: any embedding of $K_{6}$ (resp. $K_{7}$) in 3-space contains at least one link (resp. 21 links) [FM09].
Various authors have generalized the bridge index to spatial graphs, and some have shown that their version still approximates curvature as the bridge index of knots does [ST02, GY12]. There are some properties however that do not carry over to the more general setting. For instance, any knot with bridge index exactly two is prime [Sch56], but this is no longer true for $\Theta$-graphs [Mot00]. For non-prime spatial graphs, there exist estimates for the bridge indices in terms of the number of summands [TT21].
Since a random spatial graph encountered in nature usually does not have a simple diagram, it is beneficial to generalize the method of [Bla+20a] to systematically compute the bridge indices of spatial graphs from inputs that are computer-friendly. Our procedure can be summarized as follows. Given a spatial graph diagram, we pick a subset of “strands” to color and define a move that extends a color to an additional strand provided some conditions are met at a crossing. Roughly, the Wirtinger number is (a weighted count of) the minimum number of initial colorings that extend to the entire diagram by the move, minimized over all diagrams (see Section 2.3 for a formal definition). This quantity turns out to equal the bridge index, which we prove in Section 3.
@@
We use our main result, which we implement as an algorithm in Python, to provide upper estimates for the bridge index of a variety of spatial graphs. In some cases we combine our techniques with algebraic lower bounds, which we detail in Section 4, to compute exact values. We use our methods to prove in Section 5.5 that the bridge index of almost unknotted spatial graphs can be arbitrarily large. We accomplish this by modifying clasping and summing operations to be compatible with quandle colorings. Note that a connected graph $G$ with Euler characteristic $\chi(G)=\beta_{0}(G)-\beta_{1}(G)\geq 0$ is either a link or is not topologically interesting.
Theorem 1.2.
For every possible negative integer $m$, there exist almost unknotted graphs $G$ of arbitrarily large bridge index with Euler characteristic $\chi(G)=m$.
Throughout Section 5 we provide many other example computations. In particular, spatial graph whose vertices all have degree four can be obtained from link diagrams by turning some crossings into vertices, or fusing maxima and minima. We can compute exact bridge indices for spatial graphs obtained from generalized Montesinos links by fusing some maxima and minima together (see Example 5.20). In some cases, these fusing operations can still produce interesting almost unknotted graphs.
Our Python code can be found at https://github.com/hanhv/graph-wirt.
@@ Acknowledgments
Some of this work was conducted when the first author visited the other two at Arizona State University, and she thanks the mathematics department for their hospitality. We would like to thank Julien Paupert for support and helpful advice. We also thank the reviewer for many helpful suggestions that improved the paper. SB was supported by the NSF Postdoctoral Research Fellowship DMS-2303143.
@@ 2.1 Spatial graphs
A graph $G$ is a pair $(V,E)$ of a finite, non-empty set $V$ of vertices and a finite multiset $E$ of edges, where $G=E\subset V\times V$. A spatial graph is an embedding of a graph $G$ in $S^{3}$ taking vertices to points in $S^{3}$ and taking edges $(u,v)$ to an arc whose endpoints are images of $u$ and $v$. We assume that our spatial graphs are “tame” in the sense of [FH21]. We say a spatial graph is trivial if it can be isotoped to lie in the plane.
@@
A straightforward way to construct a spatial graph is to start with a link, and then add some edges. This way, the complexity of the knotted graph is obtained from the original link. This motivates the study of the following class of spatial graphs. A nontrivial spatial graph is almost unknotted [^1] Sometimes this is also referred to as minimally knotted or Brunnian in the literature. if all of its proper subgraphs are trivial. For $\Theta$-graphs, this means each constituent knot, that is, the knot obtained by deleting one of the three edges, is the unknot.
@@ 2.2 Bridge splittings
There are several definitions for the bridge index of spatial graphs, but the following formulation, which also appeared in the work of Ozawa [Oza12], is compatible with widely used concepts such as thin positions and tunnel systems. We call a properly embedded graph $\tau$ in a 3-ball $B$ a trivial tangle if components of $\tau$ are trees or arcs and there exists a disk $D$ in $B$ containing $\tau$; see Figure 1.
Figure 1: A trivial tangle, which contains some number of trees and arcs (not necessarily grouped together as shown). We did not include vertices where the tangle intersects $\partial B$, since when we glue two trivial tangles together along these vertices, we think of the resulting closed spatial graph as no longer having vertices where we glued.
A bridge splitting of a spatial graph $G$ in the $3$-sphere $S^{3}$ is a decomposition into two trivial tangles $(S^{3},G)=(B_{1},\tau_{1})\cup_{S^{2}}(B_{2},\tau_{2})$. The splitting sphere $S^{2}$ is also called the bridge sphere. For bridge splittings of links in $S^{3}$, each tangle has the same number of maxima with respect to the radial height function on the ball $B_{i}$. We call the minimum number of maxima over all possible bridge splittings the bridge index. Note however that in general the two tangles in a bridge splitting of a spatial graph may have a different number of maxima with respect to the radial height function on the ball $B_{i}$, where the graph $\tau_{i}$ is isotoped so that each component contains a single maximum. For this reason there exist multiple definitions of bridge index for spatial graphs in the literature. Our definition will be the following.
@@
The bridge index of a particular bridge splitting $(S^{3},G)=(B_{1},\tau_{1})\cup_{S^{2}}(B_{2},\tau_{2})$ of a spatial graph $G$, denoted $\beta(S^{3},G)$, is defined to be $\tfrac{1}{2}|S^{2}\cap G|$. The bridge index of a spatial graph $G$, denoted $\beta(G)$, is the minimum bridge index over all bridge splittings for $G$.
When it is clear from context, we may abuse notation and use $\beta(G)$ to denote the bridge index of a particular bridge splitting of a graph as well.
Another method for defining the bridge index for spatial graphs which has appeared previously in the literature is to count the number of components in a trivial tangle (and then minimize this quantity over all possible bridge splittings). Historically this method has been used in the context of studying $\Theta_{n}$-graphs [God97, Mot00, Tsu03], with the constraint that each tangle contains only one vertex of $G$ so that the number of components of the two tangles is the same. In general, this may not always be the case, but one possible fix is to take the minimum number of components between the two tangles.
Once the number of tangle components is known, along with the values of the weights of each component (see Definition 2.2), we can recover the number of intersection points between the spatial graph embedding and the bridge sphere, which is the sum of the weights of the components of a trivial tangle. That is, if $\min(|\tau_{1}|,|\tau_{2}|)=i$, then we have that
$\beta(S^{3},G)=\tfrac{1}{2}|S^{2}\cap G|=\tfrac{1}{2}\sum_{i}w_{i},$
where the $w_{i}$ are the weights of the $i$ components.
Furthermore, once the number of components of one of the trivial tangles $|\tau_{1}|$ is known, one can calculate the number of components of the other trivial tangle $|\tau_{2}|$ in a bridge splitting as follows. Let $\chi(G)$ be the number of vertices of $G$ minus the number of edges of $G$. To keep track of the local maxima, we put a vertex of degree two at each local maximum which is a bridge arc. At the moment, the number of vertices is $|\tau_{1}|+|\tau_{2}|.$ The number of edges can then be determined by examining the sum of the degrees of the vertices in $\tau_{1}$: $\sum_{1\leq i\leq|\tau_{1}|}d_{i}$. It follows that
@@ 2.3 Wirtinger number
Given a spatial graph $G$ in $S^{3}$, a diagram $D$ for $G$ is a generic immersion of $G$ into the plane with each crossing labeled as an overcrossing or an undercrossing. An arc of a spatial graph diagram is a subset of the diagram which starts at an undercrossing or vertex and traces the diagram until it meets another undercrossing or vertex, at which point it ends.
For the following definition, it is useful to recast a diagram as a (possibly) disconnected planar graph, where degree one vertices are placed on either side of an undercrossing and the piece of the diagram in between is deleted, as in Figure 2. For the sake of clarity, we will refer to these degree one vertices exclusively as endpoints (or free ends); that is, “vertex” will always imply degree three or more, as assumed in Section 2.1. In other words, vertices of the diagram correspond to vertices of the spatial graph, while endpoints are artifacts of the diagram which we have added.
Figure 2: When we recast a spatial graph diagram as a planar graph, we delete small neighborhoods of the undercrossings and add endpoints to each side of the deleted portion.
@@
A strand of a diagram for a spatial graph is a connected component of the diagram, when thought of as a planar graph as described above. The weight of a strand is the number of endpoints of the strand.
Arcs which both begin and end at an undercrossing, such as the green arc in Figure 3, are strands with weight $2$. Arcs which begin or end at vertices, such as the red arc in Figure 3, are part of a larger strand, but not strands themselves. Strands with only one vertex, such as the blue strand in Figure 3, have weight equal to the degree of the vertex.
Figure 3: Some examples (blue and green) and non-examples (red) of strands. The blue strand has weight $3$ and the green strand has weight $2$.
We now describe a process for coloring a spatial graph diagram, analogous to the procedure detailed in [Bla+20a] for links. First choose $k$ distinct colors, and color a subset of strands of $D$ using these $k$ colors, which we call seeds. Sometimes we will also refer to the strands which are colored the initial colors themselves as seeds. Then continue to color any uncolored arcs using the following coloring move, which extends the coloring of a colored arc which passes under another colored arc to the other arc passing under the overstrand. See Figure 4. In particular, once a vertex is reached, the coloring does not extend across the vertex; the other arcs connected to the same vertex must receive their color from a coloring move elsewhere in the diagram.
Note that the coloring move is valid both when the overstrand matches and does not match the color of the understrand; all that matters is that the overstrand has been colored. Also, while strands which are included as seeds will contain multiple arcs all of the same color (unless the strand itself is just one arc), in general arcs connected to the same vertex can end up being colored with different colors if not included as a seed. See Figure 5 for an example of this phenomenon.
@@
We say that a spatial graph diagram is completely colored if all free ends of every strand in the diagram have received a color (in particular, once all free ends of a planar tree in a diagram have been colored, we consider the entire tree to be colored). Given a spatial graph diagram, suppose it can be completely colored by starting with seeds $\{s_{1},\ldots,s_{n}\}$ and extending these colors using the coloring move. If $w_{i}$ is the weight of the seed $s_{i}$, and $\tfrac{1}{2}\sum_{i=1}^{n}w_{i}=\vcentcolon k$, then we call the diagram $k$-Wirtinger colorable. Note that we must generalize the definition in [Bla+20a] to incorporate weights into our definition in order to match the definition we have chosen for bridge index (see Definition 2.1 and the following discussion).
Example 2.3.
See Figure 5 for an example of a $\tfrac{7}{2}$-Wirtinger colorable diagram for a spatial $\Theta$-graph using three seeds. It is an exercise to check that this diagram cannot be completely colored with only two seeds, and (a slightly harder exercise to see that) furthermore, it cannot be completely colored with three seeds each of weight $2$. Thus $\tfrac{7}{2}$ is actually the smallest value of $k$ for which this diagram is $k$-Wirtinger colorable.
Figure 5: This diagram is $\tfrac{7}{2}$-Wirtinger colorable, using three seeds with weights 22 (red), 33 (blue), and 22 (green).
Example 2.4.
See Figure 6 for an example of a $3$-Wirtinger colorable diagram for a spatial $\Theta$-graph using two seeds. It is again an exercise to show that this is the smallest $k$ for which this diagram is $k$-Wirtinger colorable. In fact, our Python code checks for all possible combinations of strands given a fixed number of colors.
Figure 6: This diagram is $3$-Wirtinger colorable, using two seeds with weights 22 (red) and 44 (blue).
@@
The Wirtinger number of a spatial graph diagram $D$, denoted $\omega(D)$, is the minimal value of $k$ for which the diagram is $k$-Wirtinger colorable. The Wirtinger number of a spatial graph $G$, denoted $\omega(G)$, is the minimal Wirtinger number over all diagrams for $G$.